OpenStudy (anonymous):

The equation of line XY is (y-3) =-2/3 (x-4). What is the slope of a line perpendicular to XY?

OpenStudy (anonymous):

Select one: a. -2 b. 3 c. -2/3 d. 3/2

OpenStudy (anonymous):

@Nurali

OpenStudy (nurali):

(y-3) =-2/3 (x-4) y-3=-2/3x+8/3 y=-2/3x+8/3+3 y=-2/3x+17/3 Formula: y=mx+c m=slope so m=-2/3

OpenStudy (anonymous):

thanks